OpenStudy (anonymous):

the graph of y-4=1/4(x+3) contains which points

OpenStudy (zepp):

These two things shouldn't be equal

OpenStudy (zepp):

Assuming that the first function is \(x=y-4\) Then you the second one should be equal to x+4

OpenStudy (zepp):

grammar isn't strong todya :(

OpenStudy (zepp):

today*

OpenStudy (zepp):

@ParthKohli @Diyadiya

OpenStudy (zepp):

@Hero

hero (hero):

Why did you @ me?

OpenStudy (zepp):

And is it \[\frac{1}{4}(x+3)\] or \[\frac{1}{4(x+3)}\]

OpenStudy (zepp):

dunno, you are online lol

OpenStudy (anonymous):

(-3, 4) for \[ y-4 = \frac 1 3 (x+3) \] both sides equal 0

OpenStudy (anonymous):

@eliassaab thanks

OpenStudy (anonymous):

yw

hero (hero):

What? How is that the right one. @eliassaab , your equation is wrong?

OpenStudy (zepp):

@davonnn You didn't answer to my previous question :|

hero (hero):

It should at least be \(y = \large\frac{x + 3}{4} + 4\)

OpenStudy (anonymous):

That is the only equation that has one of the choices as solutions.

OpenStudy (uri):

i know the answer its so easy! :D

OpenStudy (anonymous):

\[ y =\large\frac{x + 3}{4} + 4\\ 4= \large\frac{-3 + 3}{4} + 4 \]

OpenStudy (anonymous):

@hero are you satisfied

hero (hero):

That's more like it

OpenStudy (anonymous):

Excellent.

OpenStudy (anonymous):

just compare with y=mx+c...